I am trying to help my daughter and i'm stuck....
refresh my memory on equations such as these..........
1. y+3 1/2=6
2. 2/3a=8
thank you

Answer by Alan3354(69443)

I am trying to help my daughter and i'm stuck....
refresh my memory on equations such as these..........
1. y+3 1/2=6
If you mean y + 3.5 = 6
Subtract 3.5 from both sides
y = 6 - 3.5
y = 2.5 or 2 1/2
--------------
2. 2/3a=8
Is it 2/(3a)? or (2/3)a?